### v1

@@ 4.1.2 Losses for reconstruction from sparse information
$\mathcal{L}_{D}$ $=\frac{1}{N_{t}}\sum_{i=0}^{N_{t}-1}\lVert{\left.\kern-1.2pt{\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t_{i}))\vphantom{\big{|}}\right|^{{\bm{\Omega}_{L}}}}-{\bm{u}}({\bm{\Omega}_{L}},t_{i})\rVert_{{\bm{\Omega}_{L}}}^{2},$ (11)
@@ 4.2 Time-windowing of the physics loss
Computing the residual of a partial differential equation is a temporal task, as shown in Eq. (1). We propose a simple time-windowing approach to allow the network to account for the temporal nature of the data. This windowing approach provides a means to compute the time-derivative $\partial_{t}{\bm{u}}$ required for the physics loss $\mathcal{L}_{P}$. The network takes time-windowed samples in as inputs, each sample consisting of $\tau$ sequential time-steps. Precisely, we first group all time-steps into non-overlapping sets of $\tau$ sequential time-steps to form the set $\mathcal{T}=\{[t_{i\tau},\dots,t_{(i+1)\tau-1}]\}_{i=0}^{\nicefrac{{N_{t}}}{{\tau}}}$. Elements of this set are treated as minibatches and are passed through the network to evaluate their output on which the physics loss is computed. The time-derivative $\partial_{t}{\bm{u}}$ is computed by applying a forward-Euler approximation across successive time-steps. Figure 2 provides an overview of the time-batching scheme and how elements of each batch are used to compute the temporal derivative.
@@
$\mathcal{L}_{P}=\frac{1}{(\tau-1)\lvert\mathcal{T}\rvert}\sum_{t\in\mathcal{T}}\sum_{i=0}^{\tau-1}\big{\lVert}\partial_{t}{\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t_{i+1}))-\mathcal{N}({\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t_{i+1}));{\bm{p}})\big{\rVert}_{{\bm{\Omega}_{H}}}^{2},$ (12)
@@
$\begin{split}\mathcal{L}_{P}=\frac{1}{(\tau-1)N_{t}}\sum_{t\in\mathcal{T}}\sum_{i=0}^{\tau-1}\big{\lVert}\partial_{t}{\bm{\eta_{\theta}}}({\bm{\Omega}},t_{i+1}))-\mathcal{N}({\bm{\eta_{\theta}}}({\bm{\zeta}}({\bm{\Omega}},t_{i+1}));{\bm{p}})\big{\rVert}_{{\bm{\Omega}}}^{2},\end{split}$ (13)
@@ 5 Data generation
$\begin{split}{\bm{u}}({\bm{\hat{\Omega}}}_{k},0)=\frac{\iota e^{2\pi i{\bm{\epsilon}}}}{\sigma\sqrt{2\pi}}e^{-\frac{1}{2}\left(\frac{\lvert{\bm{\hat{\Omega}}}_{k}\rvert}{\sigma}\right)^{2}}\\ \qquad\text{with}\quad{\bm{\epsilon}}_{i}\sim N(0,1),\end{split}$ (14)
where $\iota$ denotes the magnitude; $\sigma$ denotes the standard deviation; and ${\bm{\epsilon}}_{i}$ is a sample from a unit normal distribution. Equation (14) produces a pseudo-random field scaled by the magnitude of the wavenumber $\lvert{\bm{\hat{\Omega}}}_{k}\rvert$, where magnitude is computed pointwise for each wavenumber in the spectral grid. This scaling ensures that the resultant field has spatial structures of varying lengthscale. We take $\iota=10,\sigma=1.2$ for all simulations in order to provide an initial condition, which provides numerical stability.
As a consequence of the Nyquist-Shannon sampling criterion Canuto et al. (1988), the resolution of the spectral grid ${\bm{\hat{\Omega}}}_{k}$ places a lower bound on the spatial resolution. For a signal containing a maximal frequency $\omega_{\text{max}}$, the sampling frequency $\omega_{s}$ must satisfy the condition $\omega_{\text{max}}<\nicefrac{{\omega_{s}}}{{2}}$, therefore, we ensure that the spectral resolution satisfies $K\leq\nicefrac{{N}}{{2}}$. Violation of this condition induces spectral aliasing, in which energy content from frequencies exceeding the Nyquist limit $\nicefrac{{\omega_{s}}}{{2}}$ is fed back to the low frequencies, which amplifies energy content unphysically (Canuto et al., 1988). To prevent aliasing, in the task of uncovering solutions from biased data, we employ a spectral grid ${\bm{\hat{\Omega}}}_{k}\in\mathbb{Z}^{32\times 32}$, which corresponds to the physical grid ${\bm{\Omega}}\in\mathbb{R}^{64\times 64}$. For the task of reconstructing the solution form sparse data, we employ a high-resolution spectral grid ${\bm{\hat{\Omega}}}_{k}\in\mathbb{Z}^{35\times 35}$, which corresponds to the physical grid ${\bm{\Omega}}\in\mathbb{R}^{70\times 70}$. Approaching the Nyquist limit allows us to resolve the smallest structures possible without introducing spectral aliasing.
@@ 5.1 Physics loss in the Fourier domain
$\begin{split}\mathcal{L}_{P}=\frac{1}{(\tau-1)\lvert\mathcal{T}\rvert}\sum_{t\in\mathcal{T}}\sum_{i=0}^{\tau-1}\lVert\partial_{t}\hat{{\bm{\eta}}}_{\theta}({\bm{\zeta}}({\bm{\Omega}},t_{i+1}))-\hat{\mathcal{N}}(\hat{{\bm{\eta}}}_{\theta}({\bm{\zeta}}({\bm{\Omega}},t_{i+1})))\rVert_{\hat{\bm{\Omega}}_{k}}^{2},\end{split}$ (15)
@@
$\begin{split}\mathcal{L}_{P}=\frac{1}{(\tau-1)\lvert\mathcal{T}\rvert}\sum_{t\in\mathcal{T}}\sum_{i=0}^{\tau-1}\lVert\partial_{t}\hat{\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t_{i+1}))-\hat{\mathcal{N}}(\hat{\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t_{i+1})))\rVert_{\bm{\hat{\Omega}_{k}}}^{2},\end{split}$ (16)
@@ 6.1 Parameterisation of the bias
We parameterise the bias with a non-convex spatially varying function. We employ a modified Rastrigin parameterisation Rastrigin (1974), which is commonly used in non-convex optimization benchmarks
${\bm{\phi}}({\bm{x}};\mathcal{M},k_{\phi},u_{\text{max}})=$ (17)
$\frac{\mathcal{M}u_{\text{max}}}{2\pi^{2}+40}\left(20+\sum_{i=1}^{2}\left[({\bm{x}}_{i}-\pi)^{2}-10\cos(k_{\phi}({\bm{x}}_{i}-\pi))\right]\right)$ (18)
where $u_{\text{max}}$ is the maximum absolute velocity in the flow; $\mathcal{M}$ is the relative magnitude of the bias; $k_{\phi}$ is the wavenumber of the bias, and ${\bm{x}}_{1}$ and ${\bm{x}}_{2}$ are the streamwise and transversal coordinates respectively. Parameterising the bias with $\mathcal{M}$ and $k_{\phi}$ allows us to evaluate the performance of the methodology with respect different degrees of non-convexity of the bias.
@@ 6.2 Numerical details, data, and performance metrics
We uncover the solutions of partial differential equations from biased data for three physical systems of increasing complexity. Each systems is solved via the pseudospectral discretisation as described in Sec. 5, producing a solution by time-integration using the Euler-forward method with timestep of $\Delta t=5\times 10^{-3}$, which is chosen to satisfy of the Courant-Friedrichs-Lewy (CFL) condition to promote numerical stability. Model training is performed with $1024$ training samples and $256$ validation samples, which are pseudo-randomly selected from the time-domain of the solution. Each sample contains $\tau=2$ sequential time-steps. Following a hyperparameter study, the adam optimiser is employed Kingma and Ba (2015) with a learning rate of $3\times 10^{-4}$. The weighting factors for the loss, as described in Eq. 7, are taken as $\lambda_{P}=\lambda_{C}=10^{3}$, which are empirically determined (through a hyperparameter search) to provide stable training. Each model is trained for a total of $10^{4}$ epochs, which is chosen to provide sufficient convergence. The accuracy of prediction is quantified by the relative error on the validation dataset
$e=\sqrt{\frac{\sum_{t\in\mathcal{T}}\lVert{\bm{u}}(\bm{\Omega},t)-{\bm{\eta_{\theta}}}({\bm{\zeta}}(\bm{\Omega},t))\rVert_{\bm{\Omega}}^{2}}{\sum_{t\in\mathcal{T}}\lVert{\bm{u}}(\bm{\Omega},t)\rVert_{\bm{\Omega}}^{2}}}.$ (19)
This metric takes the magnitude of the solution into account, which allows the results from different systems to be compared. All experiments in this paper are run on a single Quadro RTX 8000 GPU.
@@ 6.3 The linear convection-diffusion equation
The linear convection-diffusion equation is used to describe transport phenomena Majda and Kramer (1999)
$\partial_{t}{\bm{u}}+{\bm{c}}\cdot\nabla{\bm{u}}={\nu}\Delta{\bm{u}},$ (20)
@@ 6.4 Nonlinear convection-diffusion equation
$\partial_{t}{\bm{u}}+\left({\bm{u}}\cdot{\nabla}\right){\bm{u}}={\nu}\Delta{\bm{u}},$ (21)
@@ 6.4.1 Two-dimensional turbulent flow
Finally, we consider a spatiotemporally chaotic flow, which is governed by the incompressible Navier-Stokes equations evaluated on a two-dimensional domain, which is also known as the Kolmogorov flow Fylladitakis (2018). The partial differential equations are expressions of the mass and momentum conservation, respectively
$\begin{split}{\nabla}\cdot{\bm{u}}&=0,\\ \partial_{t}{\bm{u}}+\left({\bm{u}}\cdot{\nabla}\right){\bm{u}}&=-{\nabla}p+{\nu}\Delta{\bm{u}}+{\bm{g}}.\end{split}$ (22)
where $p$ is the pressure field; and ${\bm{g}}$ is a body forcing, which enables the dynamics to be sustained by ensuring that the flow energy does not dissipate. The flow density is constant and assumed to be unity, i.e. the flow is incompressible. The use of a pseudospectral discretisation allows us to eliminate the pressure term and handle the continuity constraint implicitly, as shown in the spectral representation of the Navier-Stokes equations in A. The spatiotemporal dynamics are chaotic at the prescribed viscosity ${\nu}=\nicefrac{{1}}{{42}}$. The periodic forcing is ${\bm{g}}(\bm{x})=(\sin{(4\bm{x_{2}}),0})$. In order to remove the transient and focus on a statistically stationary regime, the transient time series up to $T_{t}=180.0$ is discarded.
@@ 6.5 Robustness
We analyse the robustness of the methodology by varying the wavenumber $k_{\phi}$ and magnitude $\mathcal{M}$ of the bias parameterization (17). Varying these parameters allows us to assess the degree to which the true state of a partial differential equation can be recovered when subjected to spatially varying bias with different degrees of non-convexity. To this end, we perform two parametric studies for each partial differential equation: (i) $\mathcal{M}=0.5$, $k_{\phi}\in\{1,3,5,7,9\}$; and (ii) $k_{\phi}=3$, $\mathcal{M}\in\{0.01,0.1,0.25,0.5,1.0\}$. For each case, we compute the relative error, $e$, between the predicted solution ${\bm{\eta_{\theta}}}({\bm{\zeta}}(\bm{\Omega},t))$ and the true solution ${\bm{u}}(\bm{\Omega},t)$, as defined in Eq. (19). We show the mean relative $\ell^{2}$-error for an ensemble of five computation for each test to ensure that the results are representative of the true performance. Empirically, we find that performance is robust to pseudo-random initialisation of network parameters with standard deviation of $\mathcal{O}(10^{-4})$ (result not shown). We employ the same network setups described in Sec. 6.2 with the same parameters for training. Assessing results using the same parameters for the three partial differential equations allows us to draw conclusion on the robustness of the methodology.
@@
First, in the linear convection-diffusion problem as shown in Figure 4(a), we demonstrate that the relative error is largely independent of the form of parameterised bias. The relative magnitude $\mathcal{M}$ and Rastrigin wavenumber $k_{\phi}$ have small impact on the ability of the model to uncover the solution to the partial differential equation, with the performance being consistent across all cases. The model performs best for the case $k_{\phi}=7,\mathcal{M}=0.5$, achieving a relative error of $e=2.568\times 10^{-1}$. Second, results for the nonlinear convection-diffusion case show marked improvement in comparison with the linear case, with errors for the numerical study shown in Figure 4(b). Despite the introduction of a nonlinear operator, the relative error is consistently lower regardless of the parameterisation of bias. The network remains equally capable of uncovering the true state across a wide range of modalities and magnitudes. Although the physics loss from the network predictions ${\bm{\eta_{\theta}}}({\bm{\zeta}}(\bm{\Omega},t))$ no longer scales in a linear fashion, this is beneficial for the training dynamics. This is because the second order nonlinearity promotes convexity in the loss-surface, which is then exploited by our gradient-based optimisation approach. Third, the PC-CNN is able to uncover the true state of the turbulent flow. Results shown in Figure 4(c) demonstrate an improvement in performance from the standard nonlinear convection-diffusion case. For both fixed Rastrigin wavenumber and magnitude, increasing the value of the parameter tends to decrease the relative error.

### v3

@@ 4.1.2 Losses for reconstruction from sparse information
$\mathcal{L}_{D}$ $=\frac{1}{N_{t}}\sum_{i=0}^{N_{t}-1}\lVert{\left.\kern-1.2pt{\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t_{i}))\vphantom{\big{|}}\right|^{{\bm{\Omega}_{L}}}}-{\bm{u}}({\bm{\Omega}_{L}},t_{i})\rVert_{{\bm{\Omega}_{L}}}^{2},$ (4.5)
@@ 4.2 Time-windowing of the physics loss
Computing the residual of a partial differential equation is a temporal task, as shown in Eq. (2.1). We propose a simple time-windowing approach to allow the network to account for the temporal nature of the data. This windowing approach provides a means to compute the time-derivative $\partial_{t}{\bm{u}}$ required for the physics loss $\mathcal{L}_{P}$. The network takes time-windowed samples in as inputs, each sample consisting of $\tau$ sequential time-steps. Precisely, we first group all time-steps into non-overlapping sets of $\tau$ sequential time-steps to form the set $\mathcal{T}=\{[t_{i\tau},\dots,t_{(i+1)\tau-1}]\}_{i=0}^{\nicefrac{{N_{t}}}{{\tau}}}$. Elements of this set are treated as minibatches and are passed through the network to evaluate their output on which the physics loss is computed. The time-derivative $\partial_{t}{\bm{u}}$ is computed by applying a forward-Euler approximation across successive time-steps. Figure 2 provides an overview of the time-batching scheme and how elements of each batch are used to compute the temporal derivative.
@@
$\mathcal{L}_{P}=\frac{1}{(\tau-1)\lvert\mathcal{T}\rvert}\sum_{t\in\mathcal{T}}\sum_{i=0}^{\tau-1}\big{\lVert}\partial_{t}{\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t_{i+1}))-\mathcal{N}({\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t_{i+1}));{\bm{p}})\big{\rVert}_{{\bm{\Omega}_{H}}}^{2},$ (4.6)
@@
$\begin{split}\mathcal{L}_{P}=\frac{1}{(\tau-1)N_{t}}\sum_{t\in\mathcal{T}}\sum_{i=0}^{\tau-1}\big{\lVert}\partial_{t}{\bm{\eta_{\theta}}}({\bm{\Omega}},t_{i+1}))-\mathcal{N}({\bm{\eta_{\theta}}}({\bm{\zeta}}({\bm{\Omega}},t_{i+1}));{\bm{p}})\big{\rVert}_{{\bm{\Omega}}}^{2},\end{split}$ (4.7)
@@ 5 Synthetic data generation
$\begin{split}{\bm{u}}({\bm{\hat{\Omega}}}_{k},0)=\frac{\iota e^{2\pi i{\bm{\epsilon}}}}{\sigma\sqrt{2\pi}}e^{-\frac{1}{2}\left(\frac{\lvert{\bm{\hat{\Omega}}}_{k}\rvert}{\sigma}\right)^{2}}\\ \qquad\text{with}\quad{\bm{\epsilon}}_{i}\sim N(0,1),\end{split}$ (5.1)
where $\iota$ denotes the magnitude; $\sigma$ denotes the standard deviation; and ${\bm{\epsilon}}_{i}$ is a sample from a unit normal distribution. Equation (5.1) produces a pseudo-random field scaled by the magnitude of the wavenumber $\lvert{\bm{\hat{\Omega}}}_{k}\rvert$, where magnitude is computed pointwise for each wavenumber in the spectral grid. This scaling ensures that the resultant field has spatial structures of varying lengthscale. We take $\iota=10,\sigma=1.2$ for all simulations in order to provide an initial condition, which provides numerical stability.
As a consequence of the Nyquist-Shannon sampling criterion Canuto et al. (1988), the resolution of the spectral grid ${\bm{\hat{\Omega}}}_{k}$ places a lower bound on the spatial resolution. For a signal containing a maximal frequency $\omega_{\text{max}}$, the sampling frequency $\omega_{s}$ must satisfy the condition $\omega_{\text{max}}<\nicefrac{{\omega_{s}}}{{2}}$, therefore, we ensure that the spectral resolution satisfies $K\leq\nicefrac{{N}}{{2}}$. Violation of this condition induces spectral aliasing, in which energy content from frequencies exceeding the Nyquist limit $\nicefrac{{\omega_{s}}}{{2}}$ is fed back to the low frequencies, which amplifies energy content unphysically (Canuto et al., 1988). To prevent aliasing, in the task of uncovering solutions from biased data, we employ a spectral grid ${\bm{\hat{\Omega}}}_{k}\in\mathbb{Z}^{32\times 32}$, which corresponds to the physical grid ${\bm{\Omega}}\in\mathbb{R}^{64\times 64}$. For the task of reconstructing the solution from sparse data, we employ a high-resolution spectral grid ${\bm{\hat{\Omega}}}_{k}\in\mathbb{Z}^{35\times 35}$, which corresponds to the physical grid ${\bm{\Omega}}\in\mathbb{R}^{70\times 70}$. Approaching the Nyquist limit allows us to resolve the smallest structures possible without introducing spectral aliasing.
@@ 5.1 Physics loss in the Fourier domain
$\begin{split}\mathcal{L}_{P}=\frac{1}{(\tau-1)\lvert\mathcal{T}\rvert}\sum_{t\in\mathcal{T}}\sum_{i=0}^{\tau-1}\lVert\partial_{t}\hat{{\bm{\eta}}}_{\theta}({\bm{\zeta}}({\bm{\Omega}},t_{i+1}))-\hat{\mathcal{N}}(\hat{{\bm{\eta}}}_{\theta}({\bm{\zeta}}({\bm{\Omega}},t_{i+1})))\rVert_{\hat{\bm{\Omega}}_{k}}^{2},\end{split}$ (5.2)
@@
$\begin{split}\mathcal{L}_{P}=\frac{1}{(\tau-1)\lvert\mathcal{T}\rvert}\sum_{t\in\mathcal{T}}\sum_{i=0}^{\tau-1}\lVert\partial_{t}\hat{\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t_{i+1}))-\hat{\mathcal{N}}(\hat{\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t_{i+1})))\rVert_{\bm{\hat{\Omega}_{k}}}^{2},\end{split}$ (5.3)
@@ 6.1 Parameterisation of the bias
Errors on predictions can broadly fall into two catergories: (i) aleatoric errors, which are typically referred to as noise due to environmental and sensors’ stochastic behaviour; and (ii) epistemic uncertainties, which are typically referred to as bias. We parameterise the bias with a non-convex spatially varying function to model an epistemic error. We employ a modified Rastrigin parameterisation Rastrigin (1974), which is commonly used in non-convex optimization benchmarks
${\bm{\phi}}({\bm{x}};\mathcal{M},k_{\phi},u_{\text{max}})=\frac{\mathcal{M}u_{\text{max}}}{2\pi^{2}+40}\left(20+\sum_{i=1}^{2}\left[({\bm{x}}_{i}-\pi)^{2}-10\cos(k_{\phi}({\bm{x}}_{i}-\pi))\right]\right)$ (6.1)
where $u_{\text{max}}$ is the maximum absolute velocity in the flow; $\mathcal{M}$ is the relative magnitude of the bias; $k_{\phi}$ is the wavenumber of the bias, and ${\bm{x}}_{1}$ and ${\bm{x}}_{2}$ are the streamwise and transversal coordinates respectively. Parameterising the bias with $\mathcal{M}$ and $k_{\phi}$ allows us to evaluate the performance of the methodology with respect to different degrees of non-convexity of the bias. Aleatoric noise removal is a fairly established field, and is not the focus of this task.
@@ 6.2 Numerical details, data, and performance metrics
We uncover the solutions of partial differential equations from biased data for three physical systems of increasing complexity. Each system is solved via the pseudospectral discretisation as described in Sec. 5, producing a solution by time-integration using the Euler-forward method with timestep of $\Delta t=5\times 10^{-3}$, which is chosen to satisfy of the Courant-Friedrichs-Lewy (CFL) condition to promote numerical stability. Model training is performed with $1024$ training samples and $256$ validation samples, which are pseudo-randomly selected from the time-domain of the solution. Each sample contains $\tau=2$ sequential time-steps. Following a hyperparameter study, the adam optimiser is employed Kingma & Ba (2015) with a learning rate of $3\times 10^{-4}$. The weighting factors for the loss, as described in Eq. 4.1, are taken as $\lambda_{P}=\lambda_{C}=10^{3}$, which are empirically determined (through a hyperparameter search) to provide stable training. Details on the network architecture can be found in A.1. Each model is trained for a total of $10^{4}$ epochs, which is chosen to provide sufficient convergence. The accuracy of prediction is quantified by the relative error on the validation dataset
$e=\sqrt{\frac{\sum_{t\in\mathcal{T}}\lVert{\bm{u}}(\bm{\Omega},t)-{\bm{\eta_{\theta}}}({\bm{\zeta}}(\bm{\Omega},t))\rVert_{\bm{\Omega}}^{2}}{\sum_{t\in\mathcal{T}}\lVert{\bm{u}}(\bm{\Omega},t)\rVert_{\bm{\Omega}}^{2}}}.$ (6.2)
This metric takes the magnitude of the solution into account, which allows the results from different systems to be compared. All experiments in this paper are run on a single Quadro RTX 8000 GPU. Numerical tests are run with different random initialisations (seeds) to assess the robustness of the network.
@@ 6.3 The linear convection-diffusion equation
The linear convection-diffusion equation is used to describe transport phenomena Majda & Kramer (1999)
$\partial_{t}{\bm{u}}+{\bm{c}}\cdot\nabla{\bm{u}}={\nu}\Delta{\bm{u}},$ (6.3)
@@ 6.4 Nonlinear convection-diffusion equation
$\partial_{t}{\bm{u}}+\left({\bm{u}}\cdot{\nabla}\right){\bm{u}}={\nu}\Delta{\bm{u}},$ (6.4)
@@ 6.4.1 Two-dimensional turbulent flow
Finally, we consider a spatiotemporally chaotic flow, which is governed by the incompressible Navier-Stokes equations evaluated on a two-dimensional domain with periodic boundary conditions, which is also known as the Kolmogorov flow Fylladitakis (2018). The partial differential equations are expressions of the mass and momentum conservation, respectively
$\begin{split}{\nabla}\cdot{\bm{u}}&=0,\\ \partial_{t}{\bm{u}}+\left({\bm{u}}\cdot{\nabla}\right){\bm{u}}&=-{\nabla}p+{\nu}\Delta{\bm{u}}+{\bm{g}}.\end{split}$ (6.5)
where $p$ is the pressure field; and ${\bm{g}}$ is a body forcing, which enables the dynamics to be sustained by ensuring that the flow energy does not dissipate. The flow density is constant and assumed to be unity, i.e. the flow is incompressible. The use of a pseudospectral discretisation allows us to eliminate the pressure term and handle the continuity constraint implicitly, as shown in the spectral representation of the Navier-Stokes equations in B. The spatiotemporal dynamics are chaotic at the prescribed viscosity ${\nu}=\nicefrac{{1}}{{42}}$. The periodic forcing is ${\bm{g}}(\bm{x})=(\sin{(4\bm{x_{2}}),0})$. In order to remove the transient and focus on a statistically stationary regime, the transient time series up to $T_{t}=180.0$ is discarded.
@@ 6.5 Robustness
We analyse the robustness of the methodology by varying the wavenumber $k_{\phi}$ and magnitude $\mathcal{M}$ of the bias parameterization (6.1). Varying these parameters allows us to assess the degree to which the true state of a partial differential equation can be recovered when subjected to spatially varying bias with different degrees of non-convexity. To this end, we perform two parametric studies for each partial differential equation: (i) $\mathcal{M}=0.5$, $k_{\phi}\in\{1,3,5,7,9\}$; and (ii) $k_{\phi}=3$, $\mathcal{M}\in\{0.01,0.1,0.25,0.5,1.0\}$. For each case, we compute the relative error, $e$, between the predicted solution ${\bm{\eta_{\theta}}}({\bm{\zeta}}(\bm{\Omega},t))$ and the true solution ${\bm{u}}(\bm{\Omega},t)$, as defined in Eq. (6.2). We show the mean relative $\ell^{2}$-error for an ensemble of five computation for each test to ensure that the results are representative of the true performance. Empirically, we find that performance is robust to pseudo-random initialisation of network parameters with standard deviation of $\mathcal{O}(10^{-4})$ (result not shown). We employ the same network setups described in Sec. 6.2 with the same parameters for training. Assessing results using the same parameters for the three partial differential equations allows us to draw conclusion on the robustness of the methodology.
@@
First, in the linear convection-diffusion problem as shown in Figure 4(a), we demonstrate that the relative error is largely independent of the form of parameterised bias. The relative magnitude $\mathcal{M}$ and Rastrigin wavenumber $k_{\phi}$ have small impact on the ability of the model to uncover the solution to the partial differential equation, with the performance being consistent across all cases. The model performs best for the case $k_{\phi}=7,\mathcal{M}=0.5$, achieving a relative error of $e=2.568\times 10^{-1}$. Second, results for the nonlinear convection-diffusion case show marked improvement in comparison with the linear case, with errors for the numerical study shown in Figure 4(b). Despite the introduction of a nonlinear operator, the relative error is consistently lower regardless of the parameterisation of bias. The network remains equally capable of uncovering the true state across a wide range of modalities and magnitudes. Although the physics loss from the network predictions ${\bm{\eta_{\theta}}}({\bm{\zeta}}(\bm{\Omega},t))$ no longer scales in a linear fashion, this is beneficial for the training dynamics. This is because the second order nonlinearity promotes convexity in the loss-surface, which is then exploited by our gradient-based optimisation approach. Third, the PC-CNN is able to uncover the true state of the turbulent flow. Results shown in Figure 4(c) demonstrate an improvement in performance from the standard nonlinear convection-diffusion case. For both fixed Rastrigin wavenumber and magnitude, increasing the value of the parameter tends to decrease the relative error. The relative error in Figure 4(c) decreases as the non-convexity of the bias increases. This is because it becomes increasingly simple to distinguish the bias from the underlying flow field if there is a larger discrepancy. We also emphasise that these tests were run with a fixed computational budget to ensure fair comparison – with additional training, we observe a further decrease in relative error across all experiments. (We also ran tests with biases parameterized with multiple wavenumbers. We found that our method robustly infers the biases — results not shown.)